# Quantizing graphs, one way or two?

### 19 Dec 2023 v3

@@ 3 Relations Between the Quantizations
However, while there exist energy independent scattering matrices for the Laplace (and similarly Schrödinger) operator, most $k$-independent unitary matrices do not square to the identity. So, while all $d_{v}\times d_{v}$ unitary matrices correspond to unitary vertex scattering matrices via equation (15), most $k$-independent unitary matrices are not themselves vertex scattering matrices of a graph Laplacian. This includes, for example, the DFT matrix (1). Consequently, it is natural to consider the wave propagation approach to construct a quantum graph as a distinct formulation of a quantum graph.
@@ 4 Wave Scattering of the Dirac Operator on a Graph
The absence of energy dependence in the vertex scattering matrices makes the Dirac operator on a metric graph a natural setting to establish a correspondence between a self-adjoint Hamiltonian and the wave-scattering construction [1]. Consider the zero mass Dirac operator and at each vertex let
@@
However, if we set the scattering matrix $U_{v}=\widehat{U}_{v}\otimes\mathrm{I}_{2}$ with $\widehat{U}_{v}$ a $d_{v}\times d_{v}$ unitary matrix the secular equation factorizes, as restricting $U_{v}$ to this form does not allow spin rotation at vertex scattering events and the Dirac operator decomposes into two operators acting on two component spinors. The secular equation becomes,
@@
We observe that the secular equation of a Dirac operator on a metric graph describing a particle with zero mass and where the vertex conditions do not allow spin rotations at the vertices matches the energy independent secular equation of wave propagation on the graph obtained by specifying unitary vertex scattering matrices directly. For massive particles this agreement is still obtained asymptotically in the high energy limit. In this sense the well studied wave-propagation quantization produces a subset of those quantum graphs corresponding to a self-adjoint Hamiltonian operator on a metric graph. The construction described here simplifies several similar constructions provided to address the same issue in [1]. It has the advantage of being both a physically relevant Hamiltonian, the natural restriction of the three dimensional Dirac operator to a metric graph, and being straightforward to describe with the simple requirements of zero mass and no spin rotation at the vertices.
@@ Acknowledgments
JH would like to thank Jens Marklof for raising the subject of this article at the QMATH 15 meeting and Gregory Berkolaiko for helpful discussion.

### 21 Dec 2023 v4

@@ 3 Relations Between the Quantizations
However, while there exist energy independent scattering matrices for the Laplace (and similarly Schrödinger) operator, most unitary matrices do not square to the identity. So, while all $d_{v}\times d_{v}$ unitary matrices correspond to unitary vertex scattering matrices via equation (15), most $k$-independent unitary matrices are not themselves vertex scattering matrices of a graph Laplacian. This includes, for example, the DFT matrix (1). Consequently, it is natural to consider the wave propagation approach to construct a quantum graph as a distinct formulation of a quantum graph.
@@ 4 Wave Scattering of the Dirac Operator on a Graph
The absence of energy dependence in the vertex scattering matrices makes the Dirac operator on a metric graph a natural setting to establish a correspondence between a self-adjoint Hamiltonian and the wave-scattering construction [1]. Consider the zero mass Dirac operator, and at each vertex let
@@
However, if we set the scattering matrix $U_{v}=\widehat{U}_{v}\otimes\mathrm{I}_{2}$ with $\widehat{U}_{v}$ a $d_{v}\times d_{v}$ unitary matrix, the secular equation factorizes, as restricting $U_{v}$ to this form does not allow spin rotation at vertex scattering events and the Dirac operator decomposes into two operators acting on two component spinors. The secular equation becomes,
@@
We observe that the secular equation of a Dirac operator on a metric graph, describing a particle with zero mass and where the vertex conditions do not allow spin rotations at the vertices, matches the energy independent secular equation of wave propagation on a quantum graph obtained by specifying unitary vertex scattering matrices directly. For massive particles this agreement is still obtained asymptotically in the high energy limit. In this sense the well studied wave-propagation quantization produces a subset of those quantum graphs corresponding to a self-adjoint Hamiltonian operator on a metric graph. The construction described here simplifies several similar constructions provided to address the same issue in [1]. It has the advantage of being both a physically relevant Hamiltonian, the natural restriction of the three dimensional Dirac operator to a metric graph, and being straightforward to describe with the simple requirements of zero mass and no spin rotation at the vertices.
@@ Acknowledgments
JH would like to thank Jens Marklof for raising the subject of this article at the QMATH 15 meeting and Gregory Berkolaiko, and the anonymous referees for helpful suggestions.